### v1

# EvoCoT: Overcoming the Exploration Bottleneck in Reinforcement Learning

Abstract
Reinforcement learning with verifiable reward (RLVR) has become a promising paradigm for post-training large language models (LLMs) to improve their reasoning capability. However, when the rollout accuracy is low on hard problems, the reward becomes sparse, limiting learning efficiency and causing exploration bottlenecks. Existing approaches either rely on stronger LLMs for distillation or filter out difficult problems, which limits scalability or restricts reasoning improvement through exploration.
We propose EvoCoT, a self-Evolving curriculum learning framework based on two-stage Chain-of-Thought (CoT) reasoning optimization. EvoCoT constrains the exploration space by self-generating and verifying CoT trajectories, then gradually shortens them to expand the space in a controlled way. This enables LLMs to stably learn from initially unsolved hard problems under sparse rewards. We apply EvoCoT to multiple LLM families, including Qwen, DeepSeek, and Llama. Experiments show that EvoCoT enables LLMs to solve previously unsolved problems, improves reasoning capability without external CoT supervision, and is compatible with various RL fine-tuning methods. We release the source code to support future research [^1] https://github.com/gtxygyzb/EvoCoT.
Introduction
Recently, reinforcement learning with verifiable reward (RLVR) has emerged as a promising paradigm for the post-training of large language models (LLMs). LLMs demonstrate remarkable reasoning capability in solving complex tasks, from math problems to code generation. Existing works (DeepSeek-AI et al. 2025; Liu et al. 2025b) compute rewards via rule-based verification of predicted final answers, effectively enhancing reasoning capability without relying on annotated reasoning trajectories.
Within RLVR, we expect LLMs to explore correct reasoning trajectories during rollouts to obtain rewards and gradually improve their reasoning capability (Yu et al. 2025; Shao et al. 2024). However, when the rollout accuracy is low on some hard problems, the LLM receives sparse rewards, hindering the improvement of reasoning capability. Due to the vast solution space, LLMs often face exploration bottlenecks on such problems.
In experiments, we find that LLMs often fail to sufficiently learn from hard problems, even after RLVR training. For example, when trained on GSM8K (Cobbe et al. 2021) and MATH (Hendrycks et al. 2021) training sets, Qwen2.5-7B still fails to solve 8.8% and 22.0% of the problems, respectively. These unsolved problems are still valuable for RLVR. If LLMs could exploit such problems more effectively during training, their reasoning capability could be further improved (Liu et al. 2025b).
Several recent works attempt to address this question. ❶ One category of methods depends on stronger LLMs to provide hints or reasoning trajectories for distillation (Nath et al. 2025a; Ma et al. 2025; Yan et al. 2025; Wu et al. 2025; Fu et al. 2025). For instance, LUFFY (Yan et al. 2025) mixes outputs from stronger LLMs into the GRPO candidate set and applies importance sampling to emphasize low-probability but correct actions. These methods enhance performance but require access to stronger LLMs, which is a strong assumption that imposes high costs and limits scalability, especially when training flagship models without available stronger teachers. ❷ Another category of methods attempts to control problem difficulty to facilitate curriculum learning for LLMs (Chen et al. 2025b; Bae et al. 2025; Shi et al. 2025). RLOR (Bae et al. 2025) computes the rollout accuracy for each group in a batch and retains only the problems within a predefined accuracy range. While this mitigates reward sparsity, it also filters out many hard problems that could serve as valuable training data, restricting the LLM’s reasoning improvement through exploration. A detailed comparison is provided in Table 1.
Methods ❶ Distillation-Free ❷ Unfiltered
ReLIFT (Ma et al. 2025)
AdaRFT (Shi et al. 2025)
RORL (Bae et al. 2025)
TAPO (Wu et al. 2025)
LUFFY (Yan et al. 2025)
Guide-GRPO (Nath et al. 2025a)
SRFT (Fu et al. 2025)
EvoCoT (Ours)
Table 1: The comparison between existing reinforcement learnin (RL) methods and EvoCoT.
In this paper, we aim to investigate the following question:
Key Question Can LLMs become self-evolving by effectively exploring and progressively solving harder problems, without relying on distillation from stronger LLMs?
We think that the low rollout accuracy on hard problems is primarily due to the vast solution space being far beyond the LLMs’ current reasoning capability, as shown in Figure 1. We propose EvoCoT, a self-Evolving curriculum learning framework based on two-stage Chain-of-Thought (CoT) reasoning optimization. The core idea of EvoCoT is to constrain the size of the exploration space. In Stage 1, the LLM receives problems and final answers, and generates its own CoT explanations. These CoTs are filtered and verified to construct step-by-step reasoning trajectories. In Stage 2, EvoCoT performs curriculum learning by progressively removing thinking steps from each CoT trajectory. This step-wise reduction gradually expands the exploration space in a controlled manner, increasing reasoning difficulty while enabling stable training under sparse rewards. Through self-evolving iterations, the LLM enhances its reasoning capability and generates higher-quality CoTs, progressively solving a portion of initially unsolved hard problems.
We apply EvoCoT to LLMs across diverse model families, including Qwen, DeepSeek, Llama, and DeepSeek-R1-Distill-Qwen (referred to as R1-Qwen). Experimental results demonstrate that:
  • Compared to GRPO, EvoCoT enables LLMs to overcome exploration bottlenecks on previously unsolved training-set problems, with average improvements of +4.5 for Qwen2.5-7B and +21.7 for R1-Qwen-1.5B.
  • Beyond the training set, EvoCoT transfers its learned reasoning to other math benchmarks, outperforming SimpleRL with average improvements of +2.3 on Qwen2.5-7B and +2.1 on R1-Qwen-1.5B.
  • Compared to GRPO and SFT, EvoCoT supports more effective self-exploration, achieving average improvements of +1.6 and +10.8 across all evaluated LLMs.
Related Work
Reinforcement Learning with Verifiable Reward
RLVR for LLMs has drawn considerable research attention following DeepSeek-R1 (DeepSeek-AI et al. 2025) and Kimi-k1.5 (Team et al. 2025). However, recent studies (Yue et al. 2025; Zhao et al. 2025) suggest that the performance of the RLVR-trained model is fundamentally constrained by the base model’s inherent capability, as RLVR only biases the base model’s output distribution toward reward-maximizing paths. In RLVR, rewards are sometimes too sparse compared to the large solution space, causing exploration bottlenecks that prevent finding solutions unexplored by the base model. Some works (Ma et al. 2025; Chen et al. 2025a; Fu et al. 2025; Liu et al. 2025c; Wu et al. 2025; Yan et al. 2025; Nath et al. 2025b; Goldie et al. 2025) attempt to incorporate off-policy data into training. For instance, ReLIFT (Ma et al. 2025), SASR (Chen et al. 2025a), SRFT (Fu et al. 2025) and SuperRL (Liu et al. 2025c) integrate RLVR with supervised fine-tuning (SFT). Meanwhile, TAPO (Wu et al. 2025), LUFFY (Yan et al. 2025) and Guide-GRPO (Nath et al. 2025b) leverage reference CoT or hints generated by stronger LLMs, or query an external thought library to guide policy optimization. Unfortunately, these methods either rely on distillation from stronger LLMs or high-quality training data.
Curriculum Learning for Reasoning Tasks
Curriculum learning (Bengio et al. 2009) is a training strategy that arranges examples ordered from easy to hard. Previous works (Team et al. 2025; Xie et al. 2025; Liu et al. 2025a) explore the application of curriculum learning in RLVR for LLM post-training, demonstrating that the difficulty arrangement of the RL training data is critical for achieving competitive performance. However, existing difficulty-arranging methods have some limitations. RORL (Bae et al. 2025) filters out too hard or too easy problems for the current LLM to solve, but some discarded hard problems could be valuable for training; E2H (Parashar et al. 2025), SEC (Chen et al. 2025b) and AdaRFT (Shi et al. 2025) dynamically adapt the probability distribution on difficulties for sampling, but they require fine-grained difficulty estimation in the dataset; R3 (Xi et al. 2024) and AdaBack (Amani et al. 2025) smoothly increase difficulty by showing the LLM gradually shorter prefixes of CoT, whereas they necessitate complete CoT data for training.
EvoCoT
Algorithm 1 EvoCoT: Self-Evolving Curriculum Learning
1def EvoCoT(LLM, D, T):
2 # Self-evolving Iterations
3 for t in range(T):
4 # Stage 1: Generate & Filter CoTs
5 T_set = []
6 for (Q, A) in D:
7 C = LLM.generate(Q, A) # Generate
8 $\hat{A}$ = LLM.generate(Q, C) # Verify
9 if $\hat{A}$ == A:
10 steps = split_steps(C)
11 T_set.append((Q, steps, A))
12
13 # Stage 2: Step-wise Curriculum
14 for (Q, steps, A) in T_set:
15 n = len(steps)
16 # Train from full to zero length
17 # k: retained steps count
18 for k in range(n, -1, -1):
19 $C_{k}$ = partial_CoT(Q, steps, k)
20 $C^{*},\hat{A}$ = LLM.generate($C_{k}$)
21 LLM.train(reward=($\hat{A}$ == A))
22
23 return LLM
Figure 1: The overall framework of EvoCoT. It is structured as two nested stages: Stage 1: Answer-Guided Reasoning Path Self-Generation., which generates and filters CoT trajectories from final-answer supervision, and Stage 2: Step-Wise Curriculum Learning., which implements curriculum learning by progressively shortening CoTs to increase difficulty and exploration space. The two stages iterate jointly, enabling the LLM to gradually enhance its reasoning capability through self-evolving optimization.
Self-Evolving Curriculum Learning Framework
We introduce EvoCoT, a self-evolving curriculum learning framework that enhances the LLM’s reasoning capability through iterative training on tasks of gradually increasing difficulty. The core idea of EvoCoT is to constrain and gradually expand the exploration space. As illustrated in Figure 1, EvoCoT is structured as two nested stages: the Answer-Guided Reasoning Path Self-Generation constructs CoT trajectories from final answers, and the Step-Wise Curriculum Learning implements step-wise CoT reduction for RLVR. The overall framework is outlined in Algorithm 1 and proceeds as follows:
Stage 1: Answer-Guided Reasoning Path Self-Generation. Given a training dataset consisting of questions and final answers, the LLM generates CoT trajectories that reconstruct how the answer could be derived. This stage follows the intuition that reasoning paths are easier to construct when the final answer is provided (Zelikman et al. 2022). The generated CoTs are filtered to ensure logical consistency and are organized into multi-step trajectories connecting the question to the final answer. Importantly, this stage does not require labeled CoTs or stronger LLMs, and transforms outcome-supervised data into reasoning paths in a fully self-generated manner.
Stage 2: Step-Wise Curriculum Learning. Given the reasoning paths constructed in Stage 1, Stage 2 implements the curriculum learning by progressively shortening each CoT trajectory (Xi et al. 2024). Starting from complete CoT trajectories, EvoCoT gradually removes thinking steps in reverse order, producing a series of training samples with increasing difficulty. As shown in Figure 1, shorter CoTs expand the LLM’s exploration space, making the reasoning more challenging. The step-wise reduction forms a difficulty progression, from easy samples with full guidance to hard ones requiring more exploration. Each sample is used to fine-tune the LLM with RLVR, enabling stable exploration across a range of reasoning complexities under sparse rewards.
The two stages iterate jointly, forming a self-evolving framework. The following subsections respectively introduce: ❶ how CoTs are generated and filtered in Stage 1; ❷ how curriculum learning is implemented in Stage 2 via step-wise CoT reduction; and ❸ the self-evolving iterative optimization along with the advantages of EvoCoT.
Stage 1: Answer-Guided Reasoning Path Self-Generation
In Stage 1, EvoCoT generates and filters reasoning trajectories from training data that contain only final answers. This process transforms outcome supervision into multi-step reasoning, without relying on additional human annotations or stronger teacher models.
The input consists of math problems formatted as $(Q,A)$ pairs, where $Q$ is the question and $A$ is the final answer to verify correctness. No CoT annotations or distilled data are required.
For problems that the current LLM cannot solve directly, we leverage Assumption 1 (Zelikman et al. 2022): given a question $Q$ and its final answer $A$, the LLM is more likely to generate a valid reasoning chain $C$ that explains how the answer was derived. Thus, the LLM is guided to explore and generate CoTs conditioned on $(Q,A)$:
$(Q,A)\xrightarrow[\text{LLM}]{}C$
To ensure logical consistency, we verify whether the LLM can recover the correct answer $A$ when conditioned on $(Q,C)$:
$(Q,C)\xrightarrow[\text{LLM}]{}\hat{A},\quad\text{keep }C\text{ iff }\hat{A}=A$
[…]
$(Q,C,\hat{A})\xrightarrow[\text{split}]{}(Q,\hat{C}=\{c_{1},c_{2},...,c_{n}\},\hat{A}),\quad\hat{A}=A$
where each $c_{i}$ is a clear reasoning sub-step, forming a multi-step trajectory suitable for subsequent curriculum learning. No additional constraints are applied to the self-generated $C$, allowing the LLM to explore freely. Even if the final answer $A$ is revealed in $C$, it will be reduced during Stage 2, as described next.
Stage 2: Step-Wise Curriculum Learning
In Stage 2, the LLM is trained by progressively shortening the reasoning trajectories generated in Stage 1. This process forms a curriculum that enables the LLM to perform reasoning from easy to hard within each individual sample. Training relies solely on outcome verification as the reward.
Given a complete reasoning trajectory $(Q,c_{1},c_{2},\dots,c_{n})$, training proceeds by gradually truncating the tail steps to increase difficulty. The curriculum follows:
$(Q,c_{1},c_{2},\dots,c_{n-1},c_{n})\to(C^{*},A^{*})$
$(Q,c_{1},c_{2},\dots,c_{n-1})\to(C^{*},A^{*})$
$\vdots$
$(Q,c_{1})\to(C^{*},A^{*})$
$(Q)\to(C^{*},A^{*})$
where $C^{*},A^{*}$ denotes a reasoning chain and answer generated by the LLM, without any constraint on the number or form of the subsequent thinking steps.
This design is based on Assumption 2 (Xi et al. 2024): training with longer CoT guidance is easier than training with shorter or no CoT. Starting from full-length CoTs, the LLM learns to generate correct answers under strong guidance. Gradually removing steps expands the exploration space of the LLM, increasing difficulty and encouraging the discovery of more complex reasoning paths. The step-wise curriculum within each sample stabilizes training under sparse rewards and improves the overall reasoning capability of the LLM.
Self-Evolving Iterative Optimization
EvoCoT follows a self-evolving two-stage process. In each iteration, the current LLM first generates CoT trajectories from $(Q,A)$ pairs (Stage 1). These CoTs are filtered and split into step-wise reasoning paths. Then, the LLM is trained via curriculum learning by progressively shortening the CoTs (Stage 2), increasing task difficulty. After updating the LLM’s weights, its reasoning capability improves, enabling the generation of higher-quality CoTs in the next iteration.
This self-evolving process relies on Assumption 3 (Zelikman et al. 2022): Although initial CoTs may be unsound, iterative training can improve the LLM’s reasoning capability and lead to better CoT generation. We use $\mathcal{Q}$, $\mathcal{A}$, and $\mathcal{C}$ to denote the complete datasets. The process can be represented as: The process can be represented as:
$(\mathcal{Q},\mathcal{A})$ $\xrightarrow{\mathrm{LLM}^{(t)}}\mathcal{C}^{(t)}\xrightarrow{\text{learning}}\mathrm{LLM}^{(t+1)}$
$(\mathcal{Q},\mathcal{A})$ $\xrightarrow{\mathrm{LLM}^{(t+1)}}\mathcal{C}^{(t+1)}\xrightarrow{\text{learning}}\mathrm{LLM}^{(t+2)}$
$\quad\vdots$
Note that EvoCoT is orthogonal to existing training paradigms and can be applied as a complementary stage after post-training. This orthogonality arises from its self-exploration process, which does not rely on external supervision. Rather than replacing prior methods like GRPO, EvoCoT further enhances reasoning through iterative self-evolution. EvoCoT has three main advantages:
  • Avoiding reliance on human-annotated CoTs: The LLM learns solely from automatically generated reasoning chains based on $(Q,A)$ pairs, without requiring any manual CoT labels or stronger teacher models.
  • Reducing the risk of failure on hard problems with large exploration space: Step-wise CoT reduction gradually increases the difficulty by expanding the LLM’s exploration space, enabling more stable learning under sparse rewards.
  • Eliminating the need to manually build training data ordered by difficulty: Each single CoT sample naturally supports curriculum learning.
Experiments
We conduct a large-scale experiment to evaluate EvoCoT. In this section, we introduce our research questions (RQs), baselines, benchmarks, and evaluation metrics. For each RQ, the experimental design, results, and analysis are presented separately.
Research Questions
Our experimental study is guided by the following research questions:
RQ1: Can EvoCoT solve previously unsolved training problems? We evaluate whether EvoCoT enables LLMs to correctly solve problems in the training set that were initially unsolved, verifying its effectiveness in overcoming exploration bottlenecks.
RQ2: Can EvoCoT improve generalization to unseen math problems? We evaluate whether EvoCoT enhances the LLM’s performance on a diverse set of math benchmarks that are not included in its training data.
RQ3: How effective is EvoCoT compared to other learning paradigms? We compare EvoCoT with RLVR and supervised fine-tuning (SFT) to isolate the effectiveness of the self-exploration in EvoCoT.
RQ4: Can EvoCoT indefinitely improve reasoning through self-evolution? We evaluate whether EvoCoT can continuously enhance LLM reasoning through iteration, or if the performance saturates, revealing its scalability and inherent limitations.
Experimental Setup
Baselines.
We compare EvoCoT with recent open-source RLVR works, including SimpleRL (Zeng et al. 2025), DeepScaleR (Luo et al. 2025), and Open-Reasoner-Zero (Hu et al. 2025). In addition to vanilla GRPO training, we include PRIME (Cui et al. 2025) and SEC (Chen et al. 2025b) as method-level baselines, which are recent RL improvements or curriculum learning without distillation. To ensure a fair comparison, all baselines are trained with the same data budget as EvoCoT, even when their rollouts consistently fail on hard problems.
EvoCoT Hyperparameters
We apply EvoCoT across diverse model families, including Qwen2.5-7B, Llama3.1-8B, DeepSeek-Math-7B, and DeepSeek-R1-Distill-Qwen-1.5B (referred to as R1-Qwen-1.5B). We follow the baseline models and training setup provided by DeepScaleR and SimpleRL-Zoo [^2] We use https://github.com/volcengine/verl framework for training. Other hyperparameters follow the default settings.. ❶ In Stage 1, we collect problems from the GSM8K and MATH training sets where the LLM fails to solve the problem in all 8 rollouts. For each unsolved problem, 8 reasoning paths are sampled with a temperature of 1.0. ❷ In Stage 2, we train using the hyperparameters listed in Table 2. Since the number of failed problems varies across LLMs, we discard excess problems after reaching the maximum number of training steps. All experiments are conducted on 8×A100 (40GB) GPUs.
Parameter Value Parameter Value Advantage estimator GRPO Learning rate $1\times 10^{-6}$ Train batch size 32 Mini-batch size 32 Prompt length (max) 3000 Response length (max) 5192 Samples per problem 8 Temperature 1.0 KL loss enabled Yes KL loss coefficient 0.0001 Shuffle dataset No Micro batch size 1
Table 2: EvoCoT Training Hyperparameters
Benchmarks.
We evaluate EvoCoT on a broad set of math reasoning benchmarks. Training is conducted on the train splits of GSM8K (Cobbe et al. 2021) and MATH (Hendrycks et al. 2021). For evaluation, we use the test splits of GSM8K and MATH, as well as AIME 2024, AMC 2023, Minerva Math (Lewkowycz et al. 2022), and Olympiad Bench (He et al. 2024). These benchmarks cover a wide range of mathematical domains and difficulty levels, offering a comprehensive evaluation.
Evaluation Metrics.
Following prior work (Zeng et al. 2025), we use pass@k to measure the probability that at least one correct solution is generated within $k$ attempts. In all experiments, we set $k=1$. All responses are generated with a context length of 8,192, using a decoding temperature of 0.6 and sampling 8 responses per LLM [^3] We use https://github.com/huggingface/Math-Verify and https://github.com/QwenLM/Qwen2.5-Math frameworks.. Other evaluation hyperparameters follow the default settings.
Figure 2: Number of correct rollouts over training steps during EvoCoT training.
Model GSM8K MATH Avg. Llama3.1-8B + GRPO 84.3 21.9 53.2    +EvoCoT 83.6 21.9 52.8 DeepSeek-Math-7B + GRPO 80.8 37.1 59.0    +EvoCoT 78.5 37.2 57.9 Qwen2.5-7B + GRPO 91.2 78.0 84.6    +EvoCoT 95.4 82.7 89.1 R1-Qwen-1.5B + GRPO 80.7 55.7 68.2    +EvoCoT 91.9 87.8 89.9
Table 3: Comparison on the Training Set problems before and after applying EvoCoT (Only +GRPO vs. EvoCoT).
Model GSM8K MATH AIME 24 AMC 23 Minerva Math Olympiad Bench Avg. Llama3.1-8B 39.7 13.6 0.0 2.5 4.8 3.1 10.6    + SimpleRL 78.5 23.1 0.0 5.0 4.4 6.2 19.5    + EvoCoT 80.5 23.8 0.0 7.5 4.8 5.8 20.4 DeepSeek-Math-7B 28.4 19.4 0.0 10.0 5.5 4.7 11.3    + SimpleRL 79.8 38.7 0.0 15.0 16.2 12.4 27.0    + EvoCoT 76.3 39.1 0.0 20.0 19.1 13.0 27.9 Qwen2.5-7B 88.2 64.6 3.3 30.0 25.7 30.1 40.3    + SEC [^4] Reported as-is from the original paper as lack of released code. - 76.1 17.5 51.0 - - -    + SimpleRL-Zoo 92.4 79.7 10.0 52.5 34.6 38.1 51.2    + Open-Reasoner-Zero 93.8 81.7 10.0 55.0 34.2 45.6 53.4    + PRIME (380K) 91.7 80.3 13.3 65.0 39.7 41.8 55.3    + EvoCoT 91.4 76.5 20.0 60.0 37.1 35.9 53.5 R1-Qwen-1.5B 81.1 82.8 28.8 62.9 26.5 43.3 54.2    + DeepScaleR 88.2 89.4 36.7 77.5 38.2 51.6 63.6    + EvoCoT 88.0 89.7 40.0 87.5 37.1 51.6 65.7
Table 4: Benchmark performance comparison between EvoCoT and other baselines.
RQ1: EvoCoT Overcome Exploration Bottlenecks
In RQ1, we examine whether EvoCoT enables LLMs to solve training problems that were previously unsolved. We focus on GSM8K and MATH training data, and select problems where the LLM fails to solve in rollouts. These problems are added to the EvoCoT’s training set. Figure 2 tracks the number of correct rollouts during training, while Table 3 compares performance before and after applying EvoCoT on these challenging problems.
❶ EvoCoT maintains rollout accuracy as reasoning shortens. As shown in Figure 2, most LLMs maintain the number of correct rollouts throughout training compared to GRPO, even when the length of reasoning trajectories decreases. Notably, R1-Qwen-1.5B consistently achieves over 220 correct out of 256 rollouts, showing reliable performance on initially unsolved problems. ❷ EvoCoT brings larger improvement to stronger LLMs. Table 3 shows that Qwen2.5-7B improves from 84.6 to 89.1, and R1-Qwen-1.5B improves from 68.2 to 89.9, with a remarkable +32.1 increase on MATH. In contrast, weaker LLMs like Llama3.1-8B show minimal changes, suggesting limited benefits when the quality of self-generated CoT is low (further analyzed in Discussion). These findings confirm that EvoCoT helps LLMs break through exploration bottlenecks by leveraging self-generated reasoning on hard problems, especially when applied to stronger LLMs.
RQ2: EvoCoT Improves Math Reasoning
In RQ2, we assess whether the reasoning capability learned by EvoCoT transfers to diverse math benchmarks beyond the training set. We also conduct comprehensive comparisons with all baselines.
EvoCoT consistently improves performance on math benchmarks. With EvoCoT, Qwen2.5-7B improves from 40.3 to 53.6, and R1-Qwen-1.5B improves from 54.2 to 65.7. On Olympiad Bench, R1-Qwen-1.5B achieves the highest score of 51.6. Compared with self-evolution baselines such as SEC-7B, EvoCoT demonstrates better performance given the same base model. Considering that the training data only includes GSM8K and MATH, EvoCoT ’s results are competitive with works like PRIME and Open-Reasoner that utilize broader data. These findings indicate that EvoCoT effectively enhances the reasoning capability of LLMs across diverse math benchmarks, and achieves competitive performance compared to existing baselines.
RQ3: EvoCoT Enables More Effective Self-Exploration than GRPO and SFT
Model GSM8K MATH AIME AMC Minerva Olympiad Avg. 24 23 Math Bench Llama3.1-8B 39.7 13.6 0.0 2.5 4.8 3.1 10.6 + GRPO 78.5 23.1 0.0 5.0 4.4 6.2 19.5 + SFT 61.8 20.3 0.0 10.0 7.4 7.0 17.8 + EvoCoT 80.5 23.8 0.0 7.5 4.8 5.8 20.4 DeepSeek-Math-7B 28.4 19.4 0.0 10.0 5.5 4.7 11.3 + GRPO 79.8 38.7 0.0 15.0 16.2 12.4 27.0 + SFT 46.8 25.4 0.0 2.5 4.4 6.7 14.3 + EvoCoT 76.3 39.1 0.0 20.0 19.1 13.0 27.9 Qwen2.5-7B 88.2 64.6 3.3 30.0 25.7 30.1 40.3 + GRPO 92.4 79.7 10.0 52.5 34.6 38.1 51.2 + SFT 67.9 56.7 6.7 32.5 30.5 27.3 36.9 + EvoCoT 91.4 76.5 20.0 60.0 37.1 35.9 53.5 R1-Qwen-1.5B 81.1 82.8 28.8 62.9 26.5 43.3 54.2 + GRPO 88.2 89.4 36.7 77.5 38.2 51.6 63.6 + SFT 73.6 86.6 30.0 62.5 32.0 47.4 55.3 + EvoCoT 88.0 89.7 40.0 87.5 42.8 52.0 65.7
Table 5: Performance of EvoCoT Compared to GRPO and SFT Across Different Model Families
To isolate the effectiveness of EvoCoT, we compare EvoCoT with two representative learning paradigms: RLVR implemented by GRPO, and SFT. We follow the STaR (Zelikman et al. 2022) for SFT, where each LLM generates its own CoTs, and those verified by answer consistency are used for SFT. All methods are trained on the same GSM8K and MATH datasets with equal training steps on incorrect problems. Results are shown in Table 5.
EvoCoT enables more effective self-exploration on hard problems. Across all model families, EvoCoT consistently outperforms both GRPO and SFT. On weaker LLMs such as Llama3.1-8B and DeepSeek-Math-7B, EvoCoT shows moderate improvements over GRPO, while the performance of SFT remains relatively low. On stronger LLMs, the advantage of EvoCoT becomes more noticeable. Qwen2.5-7B improves from 40.3 to 51.2 after GRPO training, and further to 53.5 with EvoCoT, where SFT is estimated to reach 36.9. R1-Qwen-1.5B reaches 65.7 with EvoCoT, exceeding 63.6 under GRPO and 55.3 under SFT. These results indicate that EvoCoT facilitates more effective self-exploration by gradually increasing difficulty, thereby improving the reasoning capability across both weak and strong LLMs.
Model GSM8K MATH AIME AMC Minerva Olympiad Avg. 24 23 Math Bench R1-Qwen-1.5B 88.2 89.4 36.7 77.5 38.2 51.6 63.6 +iteration1 87.0 89.2 36.7 80.0 40.8 52.0 64.3 +iteration2 88.0 89.7 40.0 87.5 42.8 52.0 65.7 +iteration3 89.2 90.0 40.0 87.5 36.8 51.4 65.8 Qwen2.5-7B 92.4 79.7 10.0 52.5 34.6 39.1 51.2 +iteration1 91.7 78.4 13.3 57.5 33.1 39.1 52.2 +iteration2 91.4 76.5 20.0 60.0 37.1 35.9 53.5 +iteration3 92.0 78.1 16.7 55.0 35.3 40.0 52.9 Llama3.1-8B 78.5 23.1 0.0 5.0 4.4 6.2 19.5 +iteration1 79.4 23.8 0.0 7.5 4.0 6.2 20.2 +iteration2 80.5 23.8 0.0 7.5 4.8 5.8 20.4 +iteration3 73.3 20.4 0.0 10.0 6.8 5.0 19.3
Table 6: Comparison of the Different Model Families’ Performance Across Multiple EvoCoT Iterations
RQ4: Self-Evolution Plateaus After Few Iterations
In RQ4, we investigate whether EvoCoT can continuously improve the reasoning capability of LLMs, or if the performance eventually saturates. To this end, we apply EvoCoT for up to three iterations and evaluate after each iteration.
❶ EvoCoT saturates after 1–2 iterations. As shown in Table 6, most LLMs benefit from the first or second iteration of self-evolution, but further improvements become marginal or inconsistent. For example, R1-Qwen-1.5B improves average score from 63.6 to 65.7 after two iterations, with notable increases on AMC23 (+10.0) and Minerva Math (+4.6). However, no further improvement is observed in the third iteration. A similar trend holds for Qwen2.5-7B, which increases from 51.2 to 53.5, then slightly declines to 52.9. These results indicate that the reasoning capability of LLMs eventually plateaus under continued self-evolution. ❷ Weaker LLMs exhibit early saturation. Llama3.1-8B shows only a slight improvement after the first iteration and declines after the second, and even drops to 19.3 in the third. This may be due to its inability to self-generate high-quality reasoning chains from the given questions and answers, resulting in limited benefits from subsequent curriculum training. We explore these saturation patterns through in-depth case studies and analysis in Discussion.
Discussion
Figure 3: Case study in the EvoCoT self-generated CoTs with Qwen2.5-7B. (a) A correct reasoning path. (b) Ground truth answer error in GSM8K. (c) LLM fails to generate a consistent reasoning path given (Q, A). (d) LLM forcibly splices the final answer.
In this section, we analyze why EvoCoT cannot self-evolve indefinitely. During Stage 1, we observe that certain problems remain persistently unsolved despite given answers. Representative cases are shown in Figure 3.
❶ Ground truth answer errors in the dataset. Some problems are intrinsically unlearnable due to incorrect answers in the training data. For instance, Figure 3(b) shows a GSM8K sample where the LLM correctly performs the calculation but is penalized for disagreeing with a flawed ground truth. Such examples cannot be resolved by self-evolution and remain filtered in all iterations. After manual verification, we identify over 30 such errors, accounting for roughly 10% of consistently unsolved problems.
❷ Inability to reason from (Q, A). In other cases, even when the LLM is provided with both the question and the correct answer, it fails to generate a consistent reasoning path. In Figure 3(c), the LLM rejects the provided answer and derives a different conclusion. Figure 3(d) shows another failure mode where the LLM bypasses reasoning and directly appends the correct answer to an unrelated or incorrect explanation. These reasoning paths are all filtered out before entering the curriculum learning due to inconsistency between CoTs and the given answers.
These observations lead to two key conclusions:
❶ LLMs with stronger base reasoning capabilities benefit more from EvoCoT, consistent with our experiments.
EvoCoT ultimately saturates in Stage 1: when an LLM cannot derive a valid reasoning path given (Q, A), further self-evolution is no longer possible.
Conclusion and Future Work
We present EvoCoT, a self-evolving curriculum learning framework that improves the reasoning capability of LLMs by overcoming exploration bottlenecks in RLVR. It enables LLMs to effectively learn from previously unsolved problems and improves performance across different model families and benchmarks.
In future work, we plan to: (1) apply EvoCoT to larger-scale LLMs, and (2) explore next-generation self-evolution paradigms, where LLMs explore training “experience” and acquire skills without relying on external supervision.

### v4

@@
-EvoCoT: Overcoming the Exploration Bottleneck in Reinforcement Learning
+EvoCoT: Overcoming the Exploration Bottleneck in Reinforcement Learning for LLMs
 Abstract
@@
 However, when the rollout accuracy is low on hard problems, the reward becomes sparse, limiting learning efficiency and causing exploration bottlenecks.
-Existing approaches either rely on stronger LLMs for distillation or filter out difficult problems, which limits scalability or restricts reasoning improvement through exploration.
+Existing approaches either rely on teacher models for distillation or filter out difficult problems, which limits scalability or restricts reasoning improvement through exploration.
 We propose EvoCoT, a self-Evolving curriculum learning framework based on two-stage Chain-of-Thought (CoT) reasoning optimization.
-EvoCoT constrains the exploration space by self-generating and verifying CoT trajectories, then gradually shortens them to expand the space in a controlled way.
-This enables LLMs to stably learn from initially unsolved hard problems under sparse rewards.
+EvoCoT constrains the exploration space by self-generating and verifying CoT trajectories, then gradually shortens CoT steps to expand the space in a controlled way.
+The framework enables LLMs to stably learn from initially unsolved hard problems under sparse rewards.
 We apply EvoCoT to multiple LLM families, including Qwen, DeepSeek, and Llama.
@@
 We release the source code to support future research [^1] https://github.com/gtxygyzb/EvoCoT.
-Introduction
+1 Introduction
 Recently, reinforcement learning with verifiable reward (RLVR) has emerged as a promising paradigm for the post-training of large language models (LLMs).
 LLMs demonstrate remarkable reasoning capability in solving complex tasks, from math problems to code generation.
-Existing works (DeepSeek-AI et al.
-2025; Liu et al.
-2025b) compute rewards via rule-based verification of predicted final answers, effectively enhancing reasoning capability without relying on annotated reasoning trajectories.
-Within RLVR, we expect LLMs to explore correct reasoning trajectories during rollouts to obtain rewards and gradually improve their reasoning capability (Yu et al.
-2025; Shao et al.
-2024).
+Existing works DeepSeek-AI et al.
+(2025); Liu et al.
+(2025b) compute rewards via rule-based verification of predicted final answers, effectively enhancing reasoning capability without relying on annotated reasoning trajectories.
+Within RLVR, we expect LLMs to explore correct reasoning trajectories during rollouts to obtain rewards and gradually improve their reasoning capability Yu et al.
+(2025); Shao et al.
+(2024).
 However, when the rollout accuracy is low on some hard problems, the LLM receives sparse rewards, hindering the improvement of reasoning capability.
@@
 In experiments, we find that LLMs often fail to sufficiently learn from hard problems, even after RLVR training.
-For example, when trained on GSM8K (Cobbe et al.
-2021) and MATH (Hendrycks et al.
-2021) training sets, Qwen2.5-7B still fails to solve 8.8% and 22.0% of the problems, respectively.
+For example, when trained on GSM8K Cobbe et al.
+(2021) and MATH Hendrycks et al.
+(2021) training sets, Qwen2.5-7B still fails to solve 8.8% and 22.0% of the problems, respectively (see Table 2).
 These unsolved problems are still valuable for RLVR.
-If LLMs could exploit such problems more effectively during training, their reasoning capability could be further improved (Liu et al.
-2025b).
+If LLMs could exploit such problems more effectively during training, their reasoning capability could be further improved Liu et al.
+(2025b).
 Several recent works attempt to address this question.
-❶ One category of methods depends on stronger LLMs to provide hints or reasoning trajectories for distillation (Nath et al.
-2025a; Ma et al.
-2025; Yan et al.
-2025; Wu et al.
-2025; Fu et al.
-2025).
-For instance, LUFFY (Yan et al.
-2025) mixes outputs from stronger LLMs into the GRPO candidate set and applies importance sampling to emphasize low-probability but correct actions.
-These methods enhance performance but require access to stronger LLMs, which is a strong assumption that imposes high costs and limits scalability, especially when training flagship models without available stronger teachers.
-❷ Another category of methods attempts to control problem difficulty to facilitate curriculum learning for LLMs (Chen et al.
-2025b; Bae et al.
-2025; Shi et al.
-2025).
-RLOR (Bae et al.
-2025) computes the rollout accuracy for each group in a batch and retains only the problems within a predefined accuracy range.
+❶ One category of methods depends on teacher LLMs to provide hints or reasoning trajectories for distillation Nath et al.
+(2025); Ma et al.
+(2025); Yan et al.
+(2025); Wu et al.
+(2025); Fu et al.
+(2025).
+For instance, LUFFY Yan et al.
+(2025) mixes outputs from teacher LLMs into the GRPO candidate set and applies importance sampling to emphasize low-probability but correct actions.
+These methods enhance performance but require access to teacher LLMs, which is a strong assumption that imposes high costs and limits scalability, especially when training flagship models without available teacher models.
+❷ Another category of methods attempts to control problem difficulty to facilitate curriculum learning for LLMs Chen et al.
+(2025b); Bae et al.
+(2025); Shi et al.
+(2025).
+RORL Bae et al.
+(2025) computes the rollout accuracy for each group in a batch and retains only the problems within a predefined accuracy range.
 While this mitigates reward sparsity, it also filters out many hard problems that could serve as valuable training data, restricting the LLM’s reasoning improvement through exploration.
 A detailed comparison is provided in Table 1.
+Table 1: The comparison between existing reinforcement learnin (RL) methods and EvoCoT.
 Methods ❶ Distillation-Free ❷ Unfiltered
-ReLIFT (Ma et al. 2025)
-AdaRFT (Shi et al. 2025)
-RORL (Bae et al. 2025)
-TAPO (Wu et al. 2025)
-LUFFY (Yan et al. 2025)
-Guide-GRPO (Nath et al. 2025a)
-SRFT (Fu et al. 2025)
+ReLIFT Ma et al. (2025)
+AdaRFT Shi et al. (2025)
+RORL Bae et al. (2025)
+TAPO Wu et al. (2025)
+LUFFY Yan et al. (2025)
+Guide-GRPO Nath et al. (2025)
+SRFT Fu et al. (2025)
 EvoCoT (Ours)
-Table 1: The comparison between existing reinforcement learnin (RL) methods and EvoCoT.
 In this paper, we aim to investigate the following question:
-Key Question Can LLMs become self-evolving by effectively exploring and progressively solving harder problems, without relying on distillation from stronger LLMs?
+Key Question Can LLMs become self-evolving by overcoming exploration bottlenecks and progressively enhancing reasoning capability, without distillation from teacher models?
 We think that the low rollout accuracy on hard problems is primarily due to the vast solution space being far beyond the LLMs’ current reasoning capability, as shown in Figure 1.
@@
 The core idea of EvoCoT is to constrain the size of the exploration space.
-In Stage 1, the LLM receives problems and final answers, and generates its own CoT explanations.
-These CoTs are filtered and verified to construct step-by-step reasoning trajectories.
-In Stage 2, EvoCoT performs curriculum learning by progressively removing thinking steps from each CoT trajectory.
+In Stage 1, the LLM receives problems and final answers, and generates its own CoT trajectories.
+These CoTs are filtered and verified to construct step-by-step reasoning.
+In Stage 2, EvoCoT performs curriculum learning by progressively removing reasoning steps from each CoT trajectory.
 This step-wise reduction gradually expands the exploration space in a controlled manner, increasing reasoning difficulty while enabling stable training under sparse rewards.
@@
 We apply EvoCoT to LLMs across diverse model families, including Qwen, DeepSeek, Llama, and DeepSeek-R1-Distill-Qwen (referred to as R1-Qwen). Experimental results demonstrate that:
-  • Compared to GRPO, EvoCoT enables LLMs to overcome exploration bottlenecks on previously unsolved training-set problems, with average improvements of +4.5 for Qwen2.5-7B and +21.7 for R1-Qwen-1.5B.
+  • Compared to GRPO, EvoCoT enables LLMs to overcome exploration bottlenecks on previously unsolved training set problems, with average improvements of +4.5 for Qwen2.5-7B and +21.7 for R1-Qwen-1.5B.
   • Beyond the training set, EvoCoT transfers its learned reasoning to other math benchmarks, outperforming SimpleRL with average improvements of +2.3 on Qwen2.5-7B and +2.1 on R1-Qwen-1.5B.
-  • Compared to GRPO and SFT, EvoCoT supports more effective self-exploration, achieving average improvements of +1.6 and +10.8 across all evaluated LLMs.
-Related Work
-Reinforcement Learning with Verifiable Reward
-RLVR for LLMs has drawn considerable research attention following DeepSeek-R1 (DeepSeek-AI et al.
-2025) and Kimi-k1.5 (Team et al.
-2025).
-However, recent studies (Yue et al.
-2025; Zhao et al.
-2025) suggest that the performance of the RLVR-trained model is fundamentally constrained by the base model’s inherent capability, as RLVR only biases the base model’s output distribution toward reward-maximizing paths.
+  • Compared to SFT and GRPO, EvoCoT supports more effective self-exploration, achieving average improvements of +10.8 and +1.6 across all evaluated LLMs.
+2 Related Work
+2.1 Reinforcement Learning with Verifiable Reward
+RLVR for LLMs has drawn considerable research attention following DeepSeek-R1 DeepSeek-AI et al.
+(2025) and Kimi-k1.5 Team et al.
+(2025).
+However, recent studies Yue et al.
+(2025); Zhao et al.
+(2025) suggest that the performance of the RLVR-trained model is fundamentally constrained by the base model’s inherent capability, as RLVR only biases the base model’s output distribution toward reward-maximizing paths.
 In RLVR, rewards are sometimes too sparse compared to the large solution space, causing exploration bottlenecks that prevent finding solutions unexplored by the base model.
-Some works (Ma et al.
-2025; Chen et al.
-2025a; Fu et al.
-2025; Liu et al.
-2025c; Wu et al.
-2025; Yan et al.
-2025; Nath et al.
-2025b; Goldie et al.
-2025) attempt to incorporate off-policy data into training.
-For instance, ReLIFT (Ma et al.
-2025), SASR (Chen et al.
-2025a), SRFT (Fu et al.
-2025) and SuperRL (Liu et al.
-2025c) integrate RLVR with supervised fine-tuning (SFT).
-Meanwhile, TAPO (Wu et al.
-2025), LUFFY (Yan et al.
-2025) and Guide-GRPO (Nath et al.
-2025b) leverage reference CoT or hints generated by stronger LLMs, or query an external thought library to guide policy optimization.
-Unfortunately, these methods either rely on distillation from stronger LLMs or high-quality training data.
-Curriculum Learning for Reasoning Tasks
-Curriculum learning (Bengio et al.
-2009) is a training strategy that arranges examples ordered from easy to hard.
-Previous works (Team et al.
-2025; Xie et al.
-2025; Liu et al.
-2025a) explore the application of curriculum learning in RLVR for LLM post-training, demonstrating that the difficulty arrangement of the RL training data is critical for achieving competitive performance.
+Some works Ma et al.
+(2025); Chen et al.
+(2025a); Fu et al.
+(2025); Liu et al.
+(2025c); Wu et al.
+(2025); Yan et al.
+(2025); Nath et al.
+(2025); Goldie et al.
+(2025) attempt to incorporate off-policy data into training.
+For instance, ReLIFT Ma et al.
+(2025), SASR Chen et al.
+(2025a), SRFT Fu et al.
+(2025) and SuperRL Liu et al.
+(2025c) integrate RLVR with supervised fine-tuning (SFT).
+Meanwhile, TAPO Wu et al.
+(2025), LUFFY Yan et al.
+(2025) and Guide-GRPO Nath et al.
+(2025) leverage reference CoT or hints generated by teacher models, or query an external thought library to guide policy optimization.
+Unfortunately, these methods either rely on distillation from teacher models or high-quality training data.
+2.2 Curriculum Learning for Reasoning Tasks
+Curriculum learning Bengio et al.
+(2009) is a training strategy that arranges examples ordered from easy to hard.
+In RL, curriculum learning explores strategies to balance exploration and exploitation, with methods such as promising initialization Narvekar et al.
+(2016) and reverse curriculum generation Florensa et al.
+(2017) showing effectiveness.
+However, in LLMs, overcoming exploration bottlenecks remains a major question.
+Previous works Team et al.
+(2025); Xie et al.
+(2025); Liu et al.
+(2025a) explore the application of curriculum learning in RLVR for LLM post-training, demonstrating that the difficulty arrangement of the RL training data is critical for achieving competitive performance.
 However, existing difficulty-arranging methods have some limitations.
-RORL (Bae et al.
-2025) filters out too hard or too easy problems for the current LLM to solve, but some discarded hard problems could be valuable for training; E2H (Parashar et al.
-2025), SEC (Chen et al.
-2025b) and AdaRFT (Shi et al.
-2025) dynamically adapt the probability distribution on difficulties for sampling, but they require fine-grained difficulty estimation in the dataset; R3 (Xi et al.
-2024) and AdaBack (Amani et al.
-2025) smoothly increase difficulty by showing the LLM gradually shorter prefixes of CoT, whereas they necessitate complete CoT data for training.
-EvoCoT
+RORL Bae et al.
+(2025) filters out too hard or too easy problems for the current LLM to solve, but some discarded hard problems could be valuable for training; E2H Parashar et al.
+(2025), SEC Chen et al.
+(2025b) and AdaRFT Shi et al.
+(2025) dynamically adapt the probability distribution on difficulties for sampling, but they require fine-grained difficulty estimation in the dataset; R3 Xi et al.
+(2024) and AdaBack Amani et al.
+(2025) smoothly increase difficulty by showing the LLM gradually shorter prefixes of CoT, whereas they necessitate complete CoT data for training.
+3 EvoCoT
+3.1 Self-Evolving Curriculum Learning Framework
+We introduce EvoCoT, a self-evolving curriculum learning framework for LLMs.
+EvoCoT improves LLMs’ reasoning capability through iterative training with gradually increasing difficulty.
+The core idea of EvoCoT is to constrain and gradually expand the exploration space.
+As illustrated in Figure 1, EvoCoT is structured as two nested stages: the Answer-Guided Reasoning Path Self-Generation constructs CoT trajectories from final answers, and the Step-Wise Curriculum Learning implements step-wise CoT reduction for RLVR.
+The overall pseudocode is provided in Appendix A and proceeds as follows:
+Stage 1: Answer-Guided Reasoning Path Self-Generation.
+Given a training dataset consisting of questions and final answers, the LLM generates CoT trajectories that reconstruct how the answer could be derived.
+This stage follows the intuition that reasoning paths are easier to construct when the final answer is provided.
+The generated CoTs are filtered to ensure logical consistency and are organized into multi-step trajectories connecting the question to the final answer.
+Importantly, this stage does not require annotated CoTs or teacher models, and transforms outcome-supervised data into reasoning paths in a fully self-generated manner.
+Stage 2: Step-Wise Curriculum Learning.
+Given the reasoning paths constructed in Stage 1, Stage 2 implements the curriculum learning by progressively shortening each CoT trajectory.
+Starting from complete CoT trajectories, EvoCoT gradually removes reasoning steps in reverse order, producing a series of training samples with increasing difficulty.
+As shown in Figure 1, shorter CoTs expand the LLM’s exploration space, making the reasoning more challenging.
+The step-wise reduction forms a difficulty progression, from easy samples with full guidance to hard ones requiring more exploration.
+Each sample is used to fine-tune the LLM with RLVR, enabling stable exploration across a range of reasoning complexities under sparse rewards.
+The two stages iterate jointly, forming a self-evolving framework.
+The following subsections respectively introduce: ❶ how CoTs are generated and filtered in Stage 1; ❷ how curriculum learning is implemented in Stage 2 via step-wise CoT reduction; and ❸ the self-evolving iterative optimization along with the advantages of EvoCoT.
+Figure 1: The overall framework of EvoCoT.
+It is structured as two nested stages: Stage 1: Answer-Guided Reasoning Path Self-Generation., which generates and filters CoT trajectories from final-answer supervision, and Stage 2: Step-Wise Curriculum Learning., which implements curriculum learning by progressively shortening CoTs to increase difficulty and exploration space.
+The two stages iterate jointly, enabling the LLM to gradually enhance its reasoning capability through self-evolving optimization.
+3.2 Stage 1: Answer-Guided Reasoning Path Self-Generation
+In Stage 1, EvoCoT generates and filters reasoning trajectories from training data that contain only final answers.
+This process transforms outcome supervision into multi-step reasoning, without relying on additional human annotations or teacher models.
+The training input consists of math problems formatted as $(Q,A)$ pairs, where $Q$ is the question and $A$ is the final answer.
+No CoT annotations or distilled data are required.
+For each $(Q,A)$, the LLM is prompted to generate a reasoning chain $C$ (detailed in Appendix B):
+$(Q,A)\xrightarrow[\text{LLM}]{}C$ (1)
+Intuitively, conditioning on the final answer increases the likelihood that the LLM predicts a reasoning trajectory that supports it.
+To verify consistency, we check whether the LLM can derive the correct answer $A$ when conditioned on $(Q,C)$:
+$(Q,C)\xrightarrow[\text{LLM}]{}\hat{A},\quad\text{retain }C\text{ if }\hat{A}=A.$ (2)
+[…]
+$(Q,C,\hat{A})\xrightarrow[\text{split}]{}(Q,\hat{C}=\{c_{1},c_{2},...,c_{n}\},\hat{A}),\quad\hat{A}=A$ (3)
+where each $c_{i}$ is a clear reasoning sub-step, forming a multi-step trajectory suitable for subsequent curriculum learning.
+No additional constraints are applied to the self-generated $C$, allowing the LLM to explore freely.
+3.3 Stage 2: Step-Wise Curriculum Learning
+In Stage 2, the LLM is trained by progressively shortening the reasoning trajectories generated in Stage 1.
+This process forms a curriculum that enables the LLM to perform reasoning from easy to hard within each individual sample.
+Training relies solely on outcome verification as the reward.
+Given a complete reasoning trajectory $(Q,c_{1},c_{2},\dots,c_{n})$, training proceeds by gradually truncating the tail steps to increase difficulty. The curriculum follows:
+$(Q,c_{1},c_{2},\dots,c_{n-1},c_{n})\to(C^{*},A^{*})$ (4)
+$(Q,c_{1},c_{2},\dots,c_{n-1})\to(C^{*},A^{*})$
+$\quad\quad\quad\quad\quad\quad\vdots$
+$(Q,c_{1})\to(C^{*},A^{*})$
+$(Q)\to(C^{*},A^{*})$
+where $C^{*},A^{*}$ denotes a reasoning chain and answer generated by the LLM, without any constraint on the number or form of the subsequent reasoning steps.
+Starting from full-length CoTs, the LLM learns to generate correct answers under strong guidance.
+Gradually removing steps expands the exploration space of the LLM, increasing difficulty and encouraging the discovery of more complex reasoning paths.
+The step-wise curriculum within each sample stabilizes training under sparse rewards and improves the overall reasoning capability of the LLM.
+Our design is motivated by two considerations: ❶ Training with longer CoT guidance is easier than shorter or no CoT, making the progressive reduction of steps a natural curriculum.
+❷ As trajectories shorten, the LLM needs to complement reasoning steps and ultimately derive $A$ directly from $Q$, which avoids reward hacking caused by revealing answers in the self-generated CoTs.
+3.4 Self-Evolving Iterative Optimization
+EvoCoT follows a self-evolving two-stage process.
+In each iteration, the current LLM first generates CoT trajectories from $(Q,A)$ pairs (Stage 1).
+These CoTs are filtered and split into step-wise reasoning paths.
+Then, the LLM is trained via curriculum learning by progressively shortening the CoTs (Stage 2), increasing task difficulty.
+After updating the LLM’s parameters, its reasoning capability improves, enabling the generation of higher-quality CoTs in the next iteration.
+Although initial CoTs may be imperfect, iterative training can improve the LLM’s reasoning capability and lead to better CoT generation, which in turn provides stronger guidance for subsequent learning.
+Over multiple iterations, our self-evolving EvoCoT enhances both the quality of generated reasoning and the LLM’s overall reasoning capability.
+We use $\mathcal{Q}$, $\mathcal{A}$, and $\mathcal{C}$ to denote the complete datasets.
+The $t$-th iteration can be represented as:
+$(\mathcal{Q},\mathcal{A})$ $\xrightarrow{\mathrm{LLM}^{(t)}}\mathcal{C}^{(t)}\xrightarrow{\text{learning}}\mathrm{LLM}^{(t+1)}$ (5)
+$(\mathcal{Q},\mathcal{A})$ $\xrightarrow{\mathrm{LLM}^{(t+1)}}\mathcal{C}^{(t+1)}\xrightarrow{\text{learning}}\mathrm{LLM}^{(t+2)}$
+$\quad\vdots$
+Note that EvoCoT is orthogonal to existing training paradigms and can be applied as a complementary stage after post-training.
+This orthogonality arises from its self-exploration process, which does not rely on external supervision.
+Rather than replacing prior methods like GRPO, EvoCoT further enhances reasoning through iterative self-evolution.
+EvoCoT has three main advantages:
+  • Avoiding reliance on human-annotated CoTs: The LLM learns solely from automatically generated reasoning chains based on $(Q,A)$ pairs, without requiring any manual CoT labels or teacher models.
+  • Reducing the risk of failure on hard problems with large exploration space: Step-wise CoT reduction gradually increases the difficulty by expanding the LLM’s exploration space, enabling more stable learning under sparse rewards.
+  • Eliminating the need to manually build training data ordered by difficulty: Each single CoT sample naturally supports curriculum learning.
+4 Experiments
+We conduct a large-scale experiment to evaluate EvoCoT.
+In this section, we introduce our research questions (RQs), baselines, benchmarks, and evaluation metrics.
+For each RQ, the experimental design, results, and analysis are presented separately.
+4.1 Research Questions
+Our experimental study is guided by the following research questions:
+RQ1: Can EvoCoT solve previously unsolved training problems?
+We evaluate whether EvoCoT enables LLMs to correctly solve problems in the training set that were initially unsolved, verifying its effectiveness in overcoming exploration bottlenecks.
+RQ2: Can EvoCoT improve generalization to unseen math problems?
+We evaluate whether EvoCoT enhances the LLM’s performance on a diverse set of math benchmarks that are not included in the training data.
+RQ3: How effective is EvoCoT compared to other learning paradigms? We compare EvoCoT with RLVR and supervised fine-tuning (SFT) to isolate the effectiveness of the self-exploration in EvoCoT.
+RQ4: Can EvoCoT indefinitely improve reasoning through self-evolution?
+We evaluate whether EvoCoT can continuously enhance LLM reasoning through iteration, or if the performance saturates, revealing its scalability and inherent limitations.
+4.2 Experimental Setup
+Baselines.
+We compare EvoCoT with recent open-source RLVR works, including SimpleRL Zeng et al.
+(2025), DeepScaleR Luo et al.
+(2025), and Open-Reasoner-Zero Hu et al.
+(2025).
+In addition to vanilla GRPO training, we include PRIME Cui et al.
+(2025) and SEC Chen et al.
+(2025b) as method-level baselines, which are recent RL improvements or curriculum learning without distillation.
+To ensure a fair comparison, we use the released LLMs with the prompt templates reported in the original papers, and all LLMs use the same sampling settings.
+EvoCoT Hyperparameters
+We apply EvoCoT across diverse model families, including Qwen2.5-7B Yang et al.
+(2024), Llama3.1-8B Dubey et al.
+(2024), DeepSeek-Math-7B Shao et al.
+(2024), and DeepSeek-R1-Distill-Qwen-1.5B (referred to as R1-Qwen-1.5B) DeepSeek-AI et al.
+(2025).
+We follow the baseline models and training setup provided by DeepScaleR and SimpleRL-Zoo [^2] We use https://github.com/volcengine/verl framework for training..
+❶ In Stage 1, we collect problems from the GSM8K and MATH training sets where the LLM fails to solve the problem in all 8 rollouts.
+For each unsolved problem, 8 reasoning paths are sampled with a temperature of 1.0.
+❷ In Stage 2, detailed training hyperparameters are provided in Appendix C.
+Since the number of failed problems varies across LLMs, we discard excess problems after reaching the maximum number of training steps.
+All experiments are conducted on 8×A100 (40GB) GPUs.
+Benchmarks.
+We evaluate EvoCoT on a broad set of math reasoning benchmarks.
+Training is conducted on the train splits of GSM8K Cobbe et al.
+(2021) and MATH Hendrycks et al.
+(2021).
+For evaluation, we use the test splits of GSM8K and MATH, as well as AIME 2024, AMC 2023, Minerva Math Lewkowycz et al.
+(2022), and Olympiad Bench He et al.
+(2024).
+These benchmarks cover a wide range of mathematical domains and difficulty levels, offering a comprehensive evaluation.
+Evaluation Metrics.
+Following prior work Zeng et al.
+(2025), we use pass@k to measure the probability that at least one correct solution is generated within $k$ attempts.
+In all experiments, we set $k=1$.
+All responses are generated with a context length of 8,192, using a decoding temperature of 0.6 and sampling 8 responses per LLM [^3] We use https://github.com/huggingface/Math-Verify framework..
+Other evaluation hyperparameters follow the default settings.
+Figure 2: Number of correct rollouts over training steps during EvoCoT training. Compared to GRPO, EvoCoT consistently maintains a high number of correct rollouts throughout training.
+Table 2: Comparison on the Training Set problems before and after applying EvoCoT (Only +GRPO vs. EvoCoT).
+Model GSM8K MATH Avg.
+Llama3.1-8B + GRPO 84.3 21.9 53.1
++EvoCoT 83.6 21.9 52.8
+DeepSeek-Math-7B + GRPO 80.8 37.1 59.0
++EvoCoT 78.5 37.2 57.9
+Qwen2.5-7B + GRPO 91.2 78.0 84.6
++EvoCoT 95.4 82.7 89.1
+R1-Qwen-1.5B + GRPO 80.7 55.7 68.2
++EvoCoT 91.9 87.8 89.9
+4.3 RQ1: EvoCoT Overcome Exploration Bottlenecks
+In RQ1, we examine whether EvoCoT enables LLMs to solve training problems that were previously unsolved.
+We focus on GSM8K and MATH training data, and select problems where the LLM fails to solve in rollouts.
+These problems are added to the EvoCoT’s training set.
+Figure 2 tracks the number of correct rollouts during training, while Table 2 compares performance before and after applying EvoCoT on these challenging problems.
+❶EvoCoT maintains high rollout accuracy even as reasoning shortens.
+As shown in Figure 2, EvoCoT consistently keeps correct rollouts at a high level throughout training across various LLMs, whereas GRPO drops to 0 on hard problems.
+Notably, R1-Qwen-1.5B consistently achieves over 220 correct out of 256 rollouts, showing reliable performance on initially unsolved problems.
+❷ EvoCoT brings larger improvement to stronger LLMs.
+Table 2 shows that Qwen2.5-7B improves from 84.6 to 89.1, and R1-Qwen-1.5B improves from 68.2 to 89.9, with a remarkable +32.1 increase on MATH.
+In contrast, weaker LLMs like Llama3.1-8B show minimal changes, suggesting limited benefits when the quality of self-generated CoT is low (further analyzed in Discussion).
+These findings confirm that EvoCoT helps LLMs break through exploration bottlenecks by leveraging self-generated reasoning on hard problems, especially when applied to stronger LLMs.
+4.4 RQ2: EvoCoT Generalizes LLMs’ Reasoning Capability
+Table 3: Performance comparison of EvoCoT against baselines and ablation study results.
+[…]
+In RQ2, we evaluate whether EvoCoT helps LLMs generalize reasoning capability to diverse math benchmarks beyond the training set.
+We conduct comprehensive comparisons with all baselines.
+Results are shown in Table 3.
+EvoCoT consistently improves performance on math benchmarks.
+With EvoCoT, Qwen2.5-7B improves from 40.3 to 53.5, and R1-Qwen-1.5B improves from 54.2 to 65.7.
+On Olympiad Bench, R1-Qwen-1.5B achieves the highest score of 51.6.
+Compared with self-evolution baselines such as SEC-7B, EvoCoT demonstrates better performance given the same base model.
+Considering that the training data only includes GSM8K and MATH, EvoCoT ’s results are competitive with works like PRIME and Open-Reasoner that utilize broader data.
+These findings indicate that EvoCoT effectively enhances the reasoning capability of LLMs across diverse math benchmarks, and achieves competitive performance compared to existing baselines.
+4.5 RQ3: EvoCoT Improves Self-Exploration over GRPO and SFT
+To isolate the effectiveness of EvoCoT, we conduct an ablation study comparing EvoCoT with two representative learning paradigms: RLVR implemented by GRPO, and SFT.
+Following STaR Zelikman et al.
+(2022) for SFT, each LLM generates its own CoTs, and those verified by answer consistency are used for SFT.
+All methods are trained on the same GSM8K and MATH datasets with equal training steps on incorrect problems.
+Results are shown in Table 3.
+EvoCoT enables more effective self-exploration on hard problems.
+Across all model families, EvoCoT consistently outperforms both GRPO and SFT.
+On weaker LLMs such as Llama3.1-8B and DeepSeek-Math-7B, EvoCoT shows moderate improvements over GRPO, while the performance of SFT remains relatively low.
+On stronger LLMs, the advantage of EvoCoT becomes more noticeable.
+Qwen2.5-7B improves from 40.3 to 51.2 after GRPO training, and further to 53.5 with EvoCoT, where SFT is estimated to reach 36.9.
+R1-Qwen-1.5B reaches 65.7 with EvoCoT, exceeding 63.6 under GRPO and 55.3 under SFT.
+Unlike SFT which memorizes Chu et al.
+(2025), EvoCoT gradually shortens the reasoning process and better generalizes reasoning capability.
+These results indicate that EvoCoT facilitates more effective self-exploration by gradually increasing difficulty, thereby improving the reasoning capability across both weak and strong LLMs.
+4.6 RQ4: Self-Evolution Plateaus After Few Iterations
+Table 4: Performance of Different LLM Families Across EvoCoT Iterations.
+Model GSM8K MATH AIME AMC Minerva Olympiad Avg.
+24 23 Math Bench
+R1-Qwen-1.5B 88.2 89.4 36.7 77.5 38.2 51.6 63.6
++iteration1 87.0 89.2 36.7 80.0 40.8 52.0 64.3
+[…]
++iteration3 89.2 90.0 40.0 87.5 36.8 51.4 65.8
+Qwen2.5-7B 92.4 79.7 10.0 52.5 34.6 38.1 51.2
++iteration1 91.7 78.4 13.3 57.5 33.1 39.1 52.2
+[…]
++iteration3 92.0 78.1 16.7 55.0 35.3 40.0 52.9
+Llama3.1-8B 78.5 23.1 0.0 5.0 4.4 6.2 19.5
++iteration1 79.4 23.8 0.0 7.5 4.0 6.2 20.2
+[…]
++iteration3 73.3 20.4 0.0 10.0 6.8 5.0 19.3
+In RQ4, we investigate whether EvoCoT can continuously improve the reasoning capability of LLMs, or if the performance eventually saturates.
+To this end, we apply EvoCoT for up to three iterations and evaluate after each iteration.
+❶EvoCoT saturates after 1–2 iterations.
+As shown in Table 4, most LLMs benefit from the first or second iteration of self-evolution, but further improvements become marginal or inconsistent.
+For example, R1-Qwen-1.5B improves the average score from 63.6 to 66.7 after two iterations, with notable increases on AMC23 (+10.0) and Minerva Math (+4.6).
+However, no further improvement is observed in the third iteration.
+A similar trend holds for Qwen2.5-7B, which increases from 51.2 to 53.5, then slightly declines to 52.9.
+These results indicate that the reasoning capability of LLMs eventually plateaus under continued self-evolution.
+❷ Weaker LLMs exhibit early saturation.
+Llama3.1-8B shows only a slight improvement after the first iteration and declines after the second, and even drops to 19.3 in the third.
+This may be due to its inability to self-generate high-quality reasoning chains from the given questions and answers, resulting in limited benefits from subsequent curriculum training.
+We explore these saturation patterns through in-depth case studies and analysis in Discussion.
+5 Discussion
+Figure 3: Case study in the EvoCoT self-generated CoTs with Qwen2.5-7B.
+(a) A correct reasoning path.
+(b) Ground truth answer error in GSM8K.
+(c) LLM fails to generate a consistent reasoning path given (Q, A).
+(d) LLM forcibly splices the final answer.
+In this section, we analyze why EvoCoT cannot self-evolve indefinitely.
+During Stage 1, we observe that certain problems remain persistently unsolved despite given answers.
+Representative cases are shown in Figure 3.
+❶Ground truth answer errors in the dataset.
+Some problems are intrinsically unlearnable due to incorrect answers in the training data.
+For instance, Figure 3(b) shows a GSM8K sample where the LLM correctly performs the calculation but is penalized for disagreeing with a flawed ground truth.
+Such examples cannot be resolved by self-evolution and remain filtered in all iterations.
+After manual verification, we identify over 30 such errors, accounting for roughly 10% of consistently unsolved problems.
+❷Inability to reason from (Q, A).
+In other cases, even when the LLM is provided with both the question and the correct answer, it fails to generate a consistent reasoning path.
+In Figure 3(c), the LLM rejects the provided answer and derives a different conclusion.
+Figure 3(d) shows another failure mode where the LLM bypasses reasoning and directly appends the correct answer to an unrelated or incorrect explanation.
+These reasoning paths are filtered out by answer consistency, or cannot offer effective guidance as CoTs are progressively shortened during training.
+These observations lead to two key conclusions: ❶ LLMs with stronger base reasoning capabilities benefit more from EvoCoT, consistent with our experiments.
+❷ EvoCoT ultimately saturates in Stage 1: when an LLM cannot derive a valid reasoning path given (Q, A), further self-evolution is no longer possible.
+6 Conclusion and Future Work
+We present EvoCoT, a self-evolving curriculum learning framework that improves the reasoning capability of LLMs by overcoming exploration bottlenecks in RLVR.
+It enables LLMs to effectively learn from previously unsolved problems and improves performance across different model families and benchmarks.
+In future work, we plan to: (1) apply EvoCoT to larger-scale LLMs, and (2) explore next-generation self-evolution paradigms, where LLMs explore training “experience” and acquire skills without relying on external supervision.
+Appendix
+  • Appendix A: EvoCoT Pseudocode Algorithm
+  • Appendix B: EvoCoT Prompts Templates
+  • Appendix C: EvoCoT Hyperparameters
+  • Appendix D: LLMs Usage
+Appendix A The Pseudocode of EvoCoT Algorithm
+Algorithm 1 presents the complete algorithmic workflow of EvoCoT.
 Algorithm 1 EvoCoT: Self-Evolving Curriculum Learning
@@
 23 return LLM
-Figure 1: The overall framework of EvoCoT.
-It is structured as two nested stages: Stage 1: Answer-Guided Reasoning Path Self-Generation., which generates and filters CoT trajectories from final-answer supervision, and Stage 2: Step-Wise Curriculum Learning., which implements curriculum learning by progressively shortening CoTs to increase difficulty and exploration space.
-The two stages iterate jointly, enabling the LLM to gradually enhance its reasoning capability through self-evolving optimization.
-Self-Evolving Curriculum Learning Framework
-We introduce EvoCoT, a self-evolving curriculum learning framework that enhances the LLM’s reasoning capability through iterative training on tasks of gradually increasing difficulty.
-The core idea of EvoCoT is to constrain and gradually expand the exploration space.
-As illustrated in Figure 1, EvoCoT is structured as two nested stages: the Answer-Guided Reasoning Path Self-Generation constructs CoT trajectories from final answers, and the Step-Wise Curriculum Learning implements step-wise CoT reduction for RLVR.
-The overall framework is outlined in Algorithm 1 and proceeds as follows:
-Stage 1: Answer-Guided Reasoning Path Self-Generation.
-Given a training dataset consisting of questions and final answers, the LLM generates CoT trajectories that reconstruct how the answer could be derived.
-This stage follows the intuition that reasoning paths are easier to construct when the final answer is provided (Zelikman et al.
-2022).
-The generated CoTs are filtered to ensure logical consistency and are organized into multi-step trajectories connecting the question to the final answer.
-Importantly, this stage does not require labeled CoTs or stronger LLMs, and transforms outcome-supervised data into reasoning paths in a fully self-generated manner.
-Stage 2: Step-Wise Curriculum Learning.
-Given the reasoning paths constructed in Stage 1, Stage 2 implements the curriculum learning by progressively shortening each CoT trajectory (Xi et al.
-2024).
-Starting from complete CoT trajectories, EvoCoT gradually removes thinking steps in reverse order, producing a series of training samples with increasing difficulty.
-As shown in Figure 1, shorter CoTs expand the LLM’s exploration space, making the reasoning more challenging.
-The step-wise reduction forms a difficulty progression, from easy samples with full guidance to hard ones requiring more exploration.
-Each sample is used to fine-tune the LLM with RLVR, enabling stable exploration across a range of reasoning complexities under sparse rewards.
-The two stages iterate jointly, forming a self-evolving framework.
-The following subsections respectively introduce: ❶ how CoTs are generated and filtered in Stage 1; ❷ how curriculum learning is implemented in Stage 2 via step-wise CoT reduction; and ❸ the self-evolving iterative optimization along with the advantages of EvoCoT.
-Stage 1: Answer-Guided Reasoning Path Self-Generation
-In Stage 1, EvoCoT generates and filters reasoning trajectories from training data that contain only final answers.
-This process transforms outcome supervision into multi-step reasoning, without relying on additional human annotations or stronger teacher models.
-The input consists of math problems formatted as $(Q,A)$ pairs, where $Q$ is the question and $A$ is the final answer to verify correctness. No CoT annotations or distilled data are required.
-For problems that the current LLM cannot solve directly, we leverage Assumption 1 (Zelikman et al.
-2022): given a question $Q$ and its final answer $A$, the LLM is more likely to generate a valid reasoning chain $C$ that explains how the answer was derived.
-Thus, the LLM is guided to explore and generate CoTs conditioned on $(Q,A)$:
-$(Q,A)\xrightarrow[\text{LLM}]{}C$
-To ensure logical consistency, we verify whether the LLM can recover the correct answer $A$ when conditioned on $(Q,C)$:
-$(Q,C)\xrightarrow[\text{LLM}]{}\hat{A},\quad\text{keep }C\text{ iff }\hat{A}=A$
-[…]
-$(Q,C,\hat{A})\xrightarrow[\text{split}]{}(Q,\hat{C}=\{c_{1},c_{2},...,c_{n}\},\hat{A}),\quad\hat{A}=A$
-where each $c_{i}$ is a clear reasoning sub-step, forming a multi-step trajectory suitable for subsequent curriculum learning.
-No additional constraints are applied to the self-generated $C$, allowing the LLM to explore freely.
-Even if the final answer $A$ is revealed in $C$, it will be reduced during Stage 2, as described next.
-Stage 2: Step-Wise Curriculum Learning
-In Stage 2, the LLM is trained by progressively shortening the reasoning trajectories generated in Stage 1.
-This process forms a curriculum that enables the LLM to perform reasoning from easy to hard within each individual sample.
-Training relies solely on outcome verification as the reward.
-Given a complete reasoning trajectory $(Q,c_{1},c_{2},\dots,c_{n})$, training proceeds by gradually truncating the tail steps to increase difficulty. The curriculum follows:
-$(Q,c_{1},c_{2},\dots,c_{n-1},c_{n})\to(C^{*},A^{*})$
-$(Q,c_{1},c_{2},\dots,c_{n-1})\to(C^{*},A^{*})$
-$\vdots$
-$(Q,c_{1})\to(C^{*},A^{*})$
-$(Q)\to(C^{*},A^{*})$
-where $C^{*},A^{*}$ denotes a reasoning chain and answer generated by the LLM, without any constraint on the number or form of the subsequent thinking steps.
-This design is based on Assumption 2 (Xi et al.
-2024): training with longer CoT guidance is easier than training with shorter or no CoT.
-Starting from full-length CoTs, the LLM learns to generate correct answers under strong guidance.
-Gradually removing steps expands the exploration space of the LLM, increasing difficulty and encouraging the discovery of more complex reasoning paths.
-The step-wise curriculum within each sample stabilizes training under sparse rewards and improves the overall reasoning capability of the LLM.
-Self-Evolving Iterative Optimization
-EvoCoT follows a self-evolving two-stage process.
-In each iteration, the current LLM first generates CoT trajectories from $(Q,A)$ pairs (Stage 1).
-These CoTs are filtered and split into step-wise reasoning paths.
-Then, the LLM is trained via curriculum learning by progressively shortening the CoTs (Stage 2), increasing task difficulty.
-After updating the LLM’s weights, its reasoning capability improves, enabling the generation of higher-quality CoTs in the next iteration.
-This self-evolving process relies on Assumption 3 (Zelikman et al.
-2022): Although initial CoTs may be unsound, iterative training can improve the LLM’s reasoning capability and lead to better CoT generation.
-We use $\mathcal{Q}$, $\mathcal{A}$, and $\mathcal{C}$ to denote the complete datasets.
-The process can be represented as: The process can be represented as:
-$(\mathcal{Q},\mathcal{A})$ $\xrightarrow{\mathrm{LLM}^{(t)}}\mathcal{C}^{(t)}\xrightarrow{\text{learning}}\mathrm{LLM}^{(t+1)}$
-$(\mathcal{Q},\mathcal{A})$ $\xrightarrow{\mathrm{LLM}^{(t+1)}}\mathcal{C}^{(t+1)}\xrightarrow{\text{learning}}\mathrm{LLM}^{(t+2)}$
-$\quad\vdots$
-Note that EvoCoT is orthogonal to existing training paradigms and can be applied as a complementary stage after post-training.
-This orthogonality arises from its self-exploration process, which does not rely on external supervision.
-Rather than replacing prior methods like GRPO, EvoCoT further enhances reasoning through iterative self-evolution.
-EvoCoT has three main advantages:
-  • Avoiding reliance on human-annotated CoTs: The LLM learns solely from automatically generated reasoning chains based on $(Q,A)$ pairs, without requiring any manual CoT labels or stronger teacher models.
-  • Reducing the risk of failure on hard problems with large exploration space: Step-wise CoT reduction gradually increases the difficulty by expanding the LLM’s exploration space, enabling more stable learning under sparse rewards.
-  • Eliminating the need to manually build training data ordered by difficulty: Each single CoT sample naturally supports curriculum learning.
-Experiments
-We conduct a large-scale experiment to evaluate EvoCoT.
-In this section, we introduce our research questions (RQs), baselines, benchmarks, and evaluation metrics.
-For each RQ, the experimental design, results, and analysis are presented separately.
-Research Questions
-Our experimental study is guided by the following research questions:
-RQ1: Can EvoCoT solve previously unsolved training problems?
-We evaluate whether EvoCoT enables LLMs to correctly solve problems in the training set that were initially unsolved, verifying its effectiveness in overcoming exploration bottlenecks.
-RQ2: Can EvoCoT improve generalization to unseen math problems?
-We evaluate whether EvoCoT enhances the LLM’s performance on a diverse set of math benchmarks that are not included in its training data.
-RQ3: How effective is EvoCoT compared to other learning paradigms? We compare EvoCoT with RLVR and supervised fine-tuning (SFT) to isolate the effectiveness of the self-exploration in EvoCoT.
-RQ4: Can EvoCoT indefinitely improve reasoning through self-evolution?
-We evaluate whether EvoCoT can continuously enhance LLM reasoning through iteration, or if the performance saturates, revealing its scalability and inherent limitations.
-Experimental Setup
-Baselines.
-We compare EvoCoT with recent open-source RLVR works, including SimpleRL (Zeng et al.
-2025), DeepScaleR (Luo et al.
-2025), and Open-Reasoner-Zero (Hu et al.
-2025).
-In addition to vanilla GRPO training, we include PRIME (Cui et al.
-2025) and SEC (Chen et al.
-2025b) as method-level baselines, which are recent RL improvements or curriculum learning without distillation.
-To ensure a fair comparison, all baselines are trained with the same data budget as EvoCoT, even when their rollouts consistently fail on hard problems.
-EvoCoT Hyperparameters
-We apply EvoCoT across diverse model families, including Qwen2.5-7B, Llama3.1-8B, DeepSeek-Math-7B, and DeepSeek-R1-Distill-Qwen-1.5B (referred to as R1-Qwen-1.5B).
-We follow the baseline models and training setup provided by DeepScaleR and SimpleRL-Zoo [^2] We use https://github.com/volcengine/verl framework for training.
-Other hyperparameters follow the default settings..
-❶ In Stage 1, we collect problems from the GSM8K and MATH training sets where the LLM fails to solve the problem in all 8 rollouts.
-For each unsolved problem, 8 reasoning paths are sampled with a temperature of 1.0.
-❷ In Stage 2, we train using the hyperparameters listed in Table 2.
-Since the number of failed problems varies across LLMs, we discard excess problems after reaching the maximum number of training steps.
-All experiments are conducted on 8×A100 (40GB) GPUs.
+Appendix B The Prompt Templates of EvoCoT
+This appendix provides the prompt templates used for EvoCoT Stage 1: Answer-Guided CoT Generation and Stage 2: Step-Wise Curriculum Learning.
+Figure 4 shows the Qwen2.5 prompt template.
+For other models, Stage 1 templates remain the same, while Stage 2 templates follow the special token concatenation scheme in Zeng et al.
+(2025).
+All experiments’ evaluations also use the same Stage 2 template.
+Figure 4: Qwen2.5 Prompt format used for EvoCoT
+Appendix C The Training and Evaluation Details of EvoCoT
+Table 5: EvoCoT Training Hyperparameters
 Parameter Value Parameter Value Advantage estimator GRPO Learning rate $1\times 10^{-6}$ Train batch size 32 Mini-batch size 32 Prompt length (max) 3000 Response length (max) 5192 Samples per problem 8 Temperature 1.0 KL loss enabled Yes KL loss coefficient 0.0001 Shuffle dataset No Micro batch size 1
-Table 2: EvoCoT Training Hyperparameters
-Benchmarks.
-We evaluate EvoCoT on a broad set of math reasoning benchmarks.
-Training is conducted on the train splits of GSM8K (Cobbe et al.
-2021) and MATH (Hendrycks et al.
-2021).
-For evaluation, we use the test splits of GSM8K and MATH, as well as AIME 2024, AMC 2023, Minerva Math (Lewkowycz et al.
-2022), and Olympiad Bench (He et al.
-2024).
-These benchmarks cover a wide range of mathematical domains and difficulty levels, offering a comprehensive evaluation.
-Evaluation Metrics.
-Following prior work (Zeng et al.
-2025), we use pass@k to measure the probability that at least one correct solution is generated within $k$ attempts.
-In all experiments, we set $k=1$.
-All responses are generated with a context length of 8,192, using a decoding temperature of 0.6 and sampling 8 responses per LLM [^3] We use https://github.com/huggingface/Math-Verify and https://github.com/QwenLM/Qwen2.5-Math frameworks..
-Other evaluation hyperparameters follow the default settings.
-Figure 2: Number of correct rollouts over training steps during EvoCoT training.
-Model GSM8K MATH Avg.
-Llama3.1-8B + GRPO 84.3 21.9 53.2    +EvoCoT 83.6 21.9 52.8 DeepSeek-Math-7B + GRPO 80.8 37.1 59.0    +EvoCoT 78.5 37.2 57.9 Qwen2.5-7B + GRPO 91.2 78.0 84.6    +EvoCoT 95.4 82.7 89.1 R1-Qwen-1.5B + GRPO 80.7 55.7 68.2    +EvoCoT 91.9 87.8 89.9
-Table 3: Comparison on the Training Set problems before and after applying EvoCoT (Only +GRPO vs. EvoCoT).
-Model GSM8K MATH AIME 24 AMC 23 Minerva Math Olympiad Bench Avg.
-Llama3.1-8B 39.7 13.6 0.0 2.5 4.8 3.1 10.6    + SimpleRL 78.5 23.1 0.0 5.0 4.4 6.2 19.5    + EvoCoT 80.5 23.8 0.0 7.5 4.8 5.8 20.4 DeepSeek-Math-7B 28.4 19.4 0.0 10.0 5.5 4.7 11.3    + SimpleRL 79.8 38.7 0.0 15.0 16.2 12.4 27.0    + EvoCoT 76.3 39.1 0.0 20.0 19.1 13.0 27.9 Qwen2.5-7B 88.2 64.6 3.3 30.0 25.7 30.1 40.3    + SEC [^4] Reported as-is from the original paper as lack of released code.
-- 76.1 17.5 51.0 - - -    + SimpleRL-Zoo 92.4 79.7 10.0 52.5 34.6 38.1 51.2    + Open-Reasoner-Zero 93.8 81.7 10.0 55.0 34.2 45.6 53.4    + PRIME (380K) 91.7 80.3 13.3 65.0 39.7 41.8 55.3    + EvoCoT 91.4 76.5 20.0 60.0 37.1 35.9 53.5 R1-Qwen-1.5B 81.1 82.8 28.8 62.9 26.5 43.3 54.2    + DeepScaleR 88.2 89.4 36.7 77.5 38.2 51.6 63.6    + EvoCoT 88.0 89.7 40.0 87.5 37.1 51.6 65.7
-Table 4: Benchmark performance comparison between EvoCoT and other baselines.
-RQ1: EvoCoT Overcome Exploration Bottlenecks
-In RQ1, we examine whether EvoCoT enables LLMs to solve training problems that were previously unsolved.
-We focus on GSM8K and MATH training data, and select problems where the LLM fails to solve in rollouts.
-These problems are added to the EvoCoT’s training set.
-Figure 2 tracks the number of correct rollouts during training, while Table 3 compares performance before and after applying EvoCoT on these challenging problems.
-❶ EvoCoT maintains rollout accuracy as reasoning shortens.
-As shown in Figure 2, most LLMs maintain the number of correct rollouts throughout training compared to GRPO, even when the length of reasoning trajectories decreases.
-Notably, R1-Qwen-1.5B consistently achieves over 220 correct out of 256 rollouts, showing reliable performance on initially unsolved problems.
-❷ EvoCoT brings larger improvement to stronger LLMs.
-Table 3 shows that Qwen2.5-7B improves from 84.6 to 89.1, and R1-Qwen-1.5B improves from 68.2 to 89.9, with a remarkable +32.1 increase on MATH.
-In contrast, weaker LLMs like Llama3.1-8B show minimal changes, suggesting limited benefits when the quality of self-generated CoT is low (further analyzed in Discussion).
-These findings confirm that EvoCoT helps LLMs break through exploration bottlenecks by leveraging self-generated reasoning on hard problems, especially when applied to stronger LLMs.
-RQ2: EvoCoT Improves Math Reasoning
-In RQ2, we assess whether the reasoning capability learned by EvoCoT transfers to diverse math benchmarks beyond the training set. We also conduct comprehensive comparisons with all baselines.
-EvoCoT consistently improves performance on math benchmarks.
-With EvoCoT, Qwen2.5-7B improves from 40.3 to 53.6, and R1-Qwen-1.5B improves from 54.2 to 65.7.
-On Olympiad Bench, R1-Qwen-1.5B achieves the highest score of 51.6.
-Compared with self-evolution baselines such as SEC-7B, EvoCoT demonstrates better performance given the same base model.
-Considering that the training data only includes GSM8K and MATH, EvoCoT ’s results are competitive with works like PRIME and Open-Reasoner that utilize broader data.
-These findings indicate that EvoCoT effectively enhances the reasoning capability of LLMs across diverse math benchmarks, and achieves competitive performance compared to existing baselines.
-RQ3: EvoCoT Enables More Effective Self-Exploration than GRPO and SFT
-Model GSM8K MATH AIME AMC Minerva Olympiad Avg.
-24 23 Math Bench Llama3.1-8B 39.7 13.6 0.0 2.5 4.8 3.1 10.6 + GRPO 78.5 23.1 0.0 5.0 4.4 6.2 19.5 + SFT 61.8 20.3 0.0 10.0 7.4 7.0 17.8 + EvoCoT 80.5 23.8 0.0 7.5 4.8 5.8 20.4 DeepSeek-Math-7B 28.4 19.4 0.0 10.0 5.5 4.7 11.3 + GRPO 79.8 38.7 0.0 15.0 16.2 12.4 27.0 + SFT 46.8 25.4 0.0 2.5 4.4 6.7 14.3 + EvoCoT 76.3 39.1 0.0 20.0 19.1 13.0 27.9 Qwen2.5-7B 88.2 64.6 3.3 30.0 25.7 30.1 40.3 + GRPO 92.4 79.7 10.0 52.5 34.6 38.1 51.2 + SFT 67.9 56.7 6.7 32.5 30.5 27.3 36.9 + EvoCoT 91.4 76.5 20.0 60.0 37.1 35.9 53.5 R1-Qwen-1.5B 81.1 82.8 28.8 62.9 26.5 43.3 54.2 + GRPO 88.2 89.4 36.7 77.5 38.2 51.6 63.6 + SFT 73.6 86.6 30.0 62.5 32.0 47.4 55.3 + EvoCoT 88.0 89.7 40.0 87.5 42.8 52.0 65.7
-Table 5: Performance of EvoCoT Compared to GRPO and SFT Across Different Model Families
-To isolate the effectiveness of EvoCoT, we compare EvoCoT with two representative learning paradigms: RLVR implemented by GRPO, and SFT.
-We follow the STaR (Zelikman et al.
-2022) for SFT, where each LLM generates its own CoTs, and those verified by answer consistency are used for SFT.
-All methods are trained on the same GSM8K and MATH datasets with equal training steps on incorrect problems.
-Results are shown in Table 5.
-EvoCoT enables more effective self-exploration on hard problems.
-Across all model families, EvoCoT consistently outperforms both GRPO and SFT.
-On weaker LLMs such as Llama3.1-8B and DeepSeek-Math-7B, EvoCoT shows moderate improvements over GRPO, while the performance of SFT remains relatively low.
-On stronger LLMs, the advantage of EvoCoT becomes more noticeable.
-Qwen2.5-7B improves from 40.3 to 51.2 after GRPO training, and further to 53.5 with EvoCoT, where SFT is estimated to reach 36.9.
-R1-Qwen-1.5B reaches 65.7 with EvoCoT, exceeding 63.6 under GRPO and 55.3 under SFT.
-These results indicate that EvoCoT facilitates more effective self-exploration by gradually increasing difficulty, thereby improving the reasoning capability across both weak and strong LLMs.
-Model GSM8K MATH AIME AMC Minerva Olympiad Avg.
-24 23 Math Bench R1-Qwen-1.5B 88.2 89.4 36.7 77.5 38.2 51.6 63.6 +iteration1 87.0 89.2 36.7 80.0 40.8 52.0 64.3 +iteration2 88.0 89.7 40.0 87.5 42.8 52.0 65.7 +iteration3 89.2 90.0 40.0 87.5 36.8 51.4 65.8 Qwen2.5-7B 92.4 79.7 10.0 52.5 34.6 39.1 51.2 +iteration1 91.7 78.4 13.3 57.5 33.1 39.1 52.2 +iteration2 91.4 76.5 20.0 60.0 37.1 35.9 53.5 +iteration3 92.0 78.1 16.7 55.0 35.3 40.0 52.9 Llama3.1-8B 78.5 23.1 0.0 5.0 4.4 6.2 19.5 +iteration1 79.4 23.8 0.0 7.5 4.0 6.2 20.2 +iteration2 80.5 23.8 0.0 7.5 4.8 5.8 20.4 +iteration3 73.3 20.4 0.0 10.0 6.8 5.0 19.3
-Table 6: Comparison of the Different Model Families’ Performance Across Multiple EvoCoT Iterations
-RQ4: Self-Evolution Plateaus After Few Iterations
-In RQ4, we investigate whether EvoCoT can continuously improve the reasoning capability of LLMs, or if the performance eventually saturates.
-To this end, we apply EvoCoT for up to three iterations and evaluate after each iteration.
-❶ EvoCoT saturates after 1–2 iterations.
-As shown in Table 6, most LLMs benefit from the first or second iteration of self-evolution, but further improvements become marginal or inconsistent.
-For example, R1-Qwen-1.5B improves average score from 63.6 to 65.7 after two iterations, with notable increases on AMC23 (+10.0) and Minerva Math (+4.6).
-However, no further improvement is observed in the third iteration.
-A similar trend holds for Qwen2.5-7B, which increases from 51.2 to 53.5, then slightly declines to 52.9.
-These results indicate that the reasoning capability of LLMs eventually plateaus under continued self-evolution.
-❷ Weaker LLMs exhibit early saturation.
-Llama3.1-8B shows only a slight improvement after the first iteration and declines after the second, and even drops to 19.3 in the third.
-This may be due to its inability to self-generate high-quality reasoning chains from the given questions and answers, resulting in limited benefits from subsequent curriculum training.
-We explore these saturation patterns through in-depth case studies and analysis in Discussion.
-Discussion
-Figure 3: Case study in the EvoCoT self-generated CoTs with Qwen2.5-7B.
-(a) A correct reasoning path.
-(b) Ground truth answer error in GSM8K.
-(c) LLM fails to generate a consistent reasoning path given (Q, A).
-(d) LLM forcibly splices the final answer.
-In this section, we analyze why EvoCoT cannot self-evolve indefinitely.
-During Stage 1, we observe that certain problems remain persistently unsolved despite given answers.
-Representative cases are shown in Figure 3.
-❶ Ground truth answer errors in the dataset.
-Some problems are intrinsically unlearnable due to incorrect answers in the training data.
-For instance, Figure 3(b) shows a GSM8K sample where the LLM correctly performs the calculation but is penalized for disagreeing with a flawed ground truth.
-Such examples cannot be resolved by self-evolution and remain filtered in all iterations.
-After manual verification, we identify over 30 such errors, accounting for roughly 10% of consistently unsolved problems.
-❷ Inability to reason from (Q, A).
-In other cases, even when the LLM is provided with both the question and the correct answer, it fails to generate a consistent reasoning path.
-In Figure 3(c), the LLM rejects the provided answer and derives a different conclusion.
-Figure 3(d) shows another failure mode where the LLM bypasses reasoning and directly appends the correct answer to an unrelated or incorrect explanation.
-These reasoning paths are all filtered out before entering the curriculum learning due to inconsistency between CoTs and the given answers.
-These observations lead to two key conclusions:
-❶ LLMs with stronger base reasoning capabilities benefit more from EvoCoT, consistent with our experiments.
-EvoCoT ultimately saturates in Stage 1: when an LLM cannot derive a valid reasoning path given (Q, A), further self-evolution is no longer possible.
-Conclusion and Future Work
-We present EvoCoT, a self-evolving curriculum learning framework that improves the reasoning capability of LLMs by overcoming exploration bottlenecks in RLVR.
-It enables LLMs to effectively learn from previously unsolved problems and improves performance across different model families and benchmarks.
-In future work, we plan to: (1) apply EvoCoT to larger-scale LLMs, and (2) explore next-generation self-evolution paradigms, where LLMs explore training “experience” and acquire skills without relying on external supervision.
+This appendix provides additional details on the framework and hyperparameters used for training and evaluation of EvoCoT.
+We use the Verl framework for training the models, which provides an efficient RL pipeline.
+The full list of training hyperparameters is shown in Table 5.
+For evaluation, we use the Qwen2.5-7B-Math framework [^5] https://github.com/QwenLM/Qwen2.5-Math to evaluate LLMs’ performance across various benchmarks.
+All other evaluation parameters not explicitly mentioned follow the default settings of frameworks. The specific implementation code is provided in the supplementary materials.
+Appendix D LLMs Usage
+In preparing this manuscript, we use LLMs to aid and polish the writing.
+Specifically, LLMs improve clarity, grammar, and phrasing, ensuring the text is concise and readable.
+The use of LLMs does not influence the technical contributions or the interpretation of experimental findings.
+All content polished by LLMs is carefully checked by the authors.